OpenStudy (stuck-help):

Given triangle GHI with G(4,-3),H(-4,2), and I(2,4), find the perpendicular bisector of line HI in standard form

OpenStudy (michele_laino):

we have to write the equation of the line HI, first

OpenStudy (stuck-help):

ok i dont even know how to so that

OpenStudy (michele_laino):

hint: the slope of the stright line HI, is: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{4 - 2}}{{2 - \left( { - 4} \right)}} = ...?\]

OpenStudy (michele_laino):

please write the result

OpenStudy (stuck-help):

i dont understand to even do next

OpenStudy (stuck-help):

what am i even doing do i simplify the top and bottom and then divide

OpenStudy (michele_laino):

I suppose that: \[\left( {{x_1},{y_1}} \right) = \left( { - 4,2} \right)\quad \left( {{x_2},{y_2}} \right) = \left( {2,4} \right)\] now, next step is: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{4 - 2}}{{2 - \left( { - 4} \right)}} = \frac{2}{6} = ...?\] please simplify

OpenStudy (stuck-help):

wouldnt 2-(-4) be 8 

OpenStudy (michele_laino):

we have: \(2-(-4)=6\) so the slope \(m\) is: \[\Large m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{4 - 2}}{{2 - \left( { - 4} \right)}} = \frac{2}{6} = \frac{1}{3}\]

OpenStudy (stuck-help):

i dont understand i am getting 8 for the bottom not 6

OpenStudy (michele_laino):

it is an algebraic sum, not a multiplication

OpenStudy (stuck-help):

ok

OpenStudy (stuck-help):

so the slope is 1/3

OpenStudy (stuck-help):

where do i go from there

OpenStudy (michele_laino):

from the official theory, we know that the slope \(m'\) of the line perpendicular to the line HI, is: \[\Large m' = - \frac{1}{m} = - \frac{1}{{\left( {\frac{1}{3}} \right)}} = ...?\] please continue

OpenStudy (stuck-help):

i really dont understand if you can explain how i do it then i can try

OpenStudy (michele_laino):

it is simple, you have to multiply the first fraction for the inverse of the second fraction: \[\Large m' = - \frac{1}{m} = - \frac{1}{{\left( {\frac{1}{3}} \right)}} = - \frac{1}{1} \times \frac{3}{1} = ...?\]

OpenStudy (stuck-help):

\[\frac{ -3 }{ -1 }\]

OpenStudy (michele_laino):

we have only \(one\) minus sign

OpenStudy (michele_laino):

so?

OpenStudy (michele_laino):

please retry

OpenStudy (michele_laino):

hint: \[\Large m' = - \frac{1}{m} = - \frac{1}{{\left( {\frac{1}{3}} \right)}} = - \frac{1}{1} \times \frac{3}{1} = - 3\] am I right?

OpenStudy (stuck-help):

\[\frac{ -3 }{ 1 }\]

OpenStudy (stuck-help):

yes you are right i just wrote mine in fraction

OpenStudy (michele_laino):

yes! bthat's right! Please think about the point \(G=(4,-3)\) the perpendicular line with respect to the line \(HI\), which passes at point \(G\), is given by the subsequent equation: \[\Large y - \left( { - 3} \right) = - 3\left( {x - 4} \right)\]

OpenStudy (michele_laino):

oops.. that's*

OpenStudy (stuck-help):

sorry internet is slow

OpenStudy (michele_laino):

ok! no problem. We have to simplify such equation. Here is the next step: \[\Large y + 3 = - 3x + 12\]

OpenStudy (stuck-help):

ok so \[y+3=-3x+12\] to\[y=-3x+15\] ?

OpenStudy (stuck-help):

is that right

OpenStudy (michele_laino):

please, we have to subtract\(3\) to both sides, so we get: \[\Large y + 3 - 3 = - 3x + 12 - 3\]

OpenStudy (stuck-help):

oh yeah i dont know what i was thinking \[y=-3x+9\]

OpenStudy (michele_laino):

that's right!

OpenStudy (stuck-help):

ok is there anything else

OpenStudy (michele_laino):

please note that, such line is perpendicular to the line \(HI\), nevertheless it doesn't pass at mid point of the segment \(HI\)

OpenStudy (stuck-help):

ok thank you so much

OpenStudy (michele_laino):

furthermore, if we add \(3x-9\) to both sides, we can write: \[\Large 3x + y - 9 = 0\] which is another way to write such perpendicular line